# FARM: Functional Group-Aware Representations for Small Molecules

### v3

@@ 3.2 Atom-Level Representation Learning
We employ a masked language model architecture that takes FG-enhanced SMILES as input, using atom-level tokenization and leveraging masked atom prediction as a self-supervised task to train the model. Specifically, we adopt the BERT (Devlin, 2018) with the following loss:
@@
where $\mathcal{M}$ denotes the set of indices corresponding to the masked tokens, $x_{i}$ is the original token at position $i$, and $x_{\setminus i}$ represents the input sequence with the token at position $i$ masked. To evaluate the impact of masking, we conduct experiments with varying masking percentages and test the model’s performance on six downstream tasks. The results indicate that a masking percentage of 35% yields the highest average performance across tasks. Detailed results are provided in the Appendix D.1.
@@ 3.3 Molecular Structure Learning
To learn the structural view of molecules, we employ a two-step process. First, we derive FG embeddings from the FG knowledge graph. Next, we use these embeddings as inputs for a link prediction model to predict interactions between FGs. This approach effectively captures the relationships among functional groups, both in terms of their structural characteristics and interactions, thereby facilitating the implicit learning of molecular structure.
@@
where $E^{+}$ denotes the set of positive triples, $E^{-}$ denotes the set of negative triples, and $\gamma$ represents the margin. Optimizing this loss function will minimize the score of positive triples $(\mathbf{h},\mathbf{r},\mathbf{t})$, while maximizing the score of negative triples $(\mathbf{h^{\prime}},\mathbf{r},\mathbf{t^{\prime}})$, with a margin $\gamma$ separating the two. This is achieved through margin-based ranking loss, where the function $f(\mathbf{h},\mathbf{r},\mathbf{t})$ evaluates the plausibility of the triples. The goal is to ensure that the score for positive triples is higher than that for negative triples by at least the margin $\gamma$, thus pushing positive triples closer and negative triples farther apart in the embedding space. The detailed implementation of ComplEx is described in Appendix D.2.
@@
Link prediction. For link prediction with a graph convolutional network (GCN), we first segment molecules into functional groups through FG-aware molecular segmentation, connecting each group via single bonds. We then utilize embeddings from the FG knowledge graph as node features for the GCN. The training involves computing node embeddings through graph convolution:
$\mathbf{h}_{i}^{\prime}=\text{ReLU}\left(\mathbf{W}\cdot\frac{1}{|\mathcal{N}(i)|}\sum_{j\in\mathcal{N}(i)}\mathbf{h}_{j}\right)$
where $\mathbf{h}_{i}^{\prime}$ is the updated embedding for node $i$, computed by averaging the embeddings $\mathbf{h}_{j}$ of neighboring nodes $\mathcal{N}(i)$, applying the weight matrix $\mathbf{W}$, and passing through ReLU activation function. The score estimates the probability of connections between nodes is computed with a multi-layer perceptron (MLP) and a sigmoid function: $p_{ij}=\sigma(\text{MLP}(\mathbf{h}_{i}\oplus\mathbf{h}_{j}))$.
We then sample positive edges $E^{+}$ and negative edges $E^{-}$, optimizing the model to maximize scores for positive edges and minimize those for negative ones using the loss function:
@@ 3.4 Molecular Structure Integration via Contrastive Learning
To integrate FG-enhanced SMILES representations with molecular structure information, we employ contrastive learning to align these representations with FG graph embeddings. This method captures both the atom-level and topological aspects of molecular structures, allowing the model to develop a unified representation that integrates chemical context with overall molecular architecture.
@@
Specifically, given a positive pair $(\mathbf{h}_{\text{MLM}},\mathbf{h}_{\text{pos}})$, where $\mathbf{h}_{\text{MLM}}$ is the atom-level representation derived from a pretrained BERT model and $\mathbf{h}_{\text{pos}}$ is the corresponding structure representation from a graph neural network (GNN), and a negative pair $(\mathbf{h}_{\text{MLM}},\mathbf{h}_{\text{neg}})$, where $\mathbf{h}_{\text{neg}}$ is a augmented FG-graph, the contrastive loss can be written as: $\mathcal{L}_{\text{CL}}=\frac{1}{N}\sum_{i=1}^{N}\max\left(0,\gamma-\text{cos}(\mathbf{h}_{\text{MLM}},\mathbf{h}_{\text{pos}})+\text{cos}(\mathbf{h}_{\text{MLM}},\mathbf{h}_{\text{neg}})\right)$.
@@ 4 Pre-training Data Collection and Diversity Assessment
Table 1: Evaluation of the FG-enhanced SMILES lexicon size across various databases.
[…]
ZINC15 3M 10 [5; 63] 1,151 1,089
ChEMBL25 1.8M 35 [1; 867] 15,269 10,016
Collected dataset 20M 46 [1; 867] 22,364 14,741
Table 2: Performance comparison of FARM and baseline models on MoleculeNet classification tasks. The first 10 models incorporate FG (functional group) information to enhance representation learning.
Physiology Biophysics
Dataset BBBP Tox21 ToxCast SIDER ClinTox BACE MUV HIV
#tasks 1 12 617 27 2 1 17 1
#samples 2,039 7,831 8,575 1,427 1,478 1,513 93,807 41,127
Evaluation Metric ROC-AUC (%) ($\uparrow$) ROC-AUC (%) ($\uparrow$)
MICRO (Zhang et al., 2020) 84.4 $\pm$ 1.1 77.0 $\pm$ 0.8 65.2 $\pm$ 0.8 56.7 $\pm$ 0.9 77.0 $\pm$ 2.0 77.2 $\pm$ 2.0 - 75.1 $\pm$ 1.1
MGSSL (Zhang et al., 2021) 69.7 $\pm$ 0.9 76.5 $\pm$ 0.3 64.1 $\pm$ 0.7 61.8 $\pm$ 0.8 80.7 $\pm$ 2.1 79.1 $\pm$ 0.9 78.7 $\pm$ 1.5 78.8 $\pm$ 1.2
MoleOOD (Yang et al., 2022) 71.0 $\pm$ 0.8 - - 63.4 $\pm$ 0.7 - 84.3 $\pm$ 1.1 - 79.4 $\pm$ 0.5
MCM (Wang et al., 2023) 90.0 $\pm$ 3.1 80.2 $\pm$ 1.5 - 62.7 $\pm$ 2.8 65.5 $\pm$ 1.4 82.0 $\pm$ 5.5 - -
HimGNN (Han et al., 2023) 92.8 $\pm$ 2.7 80.7 $\pm$ 1.7 - 64.2 $\pm$ 2.3 91.7 $\pm$ 3.0 85.6 $\pm$ 3.4 - -
FG-BERT (Li et al., 2023) 70.2 $\pm$ 0.9 78.4 $\pm$ 0.8 63.3 $\pm$ 0.8 64.0 $\pm$ 0.7 83.2 $\pm$ 1.6 84.5 $\pm$ 1.5 75.3 $\pm$ 2.4 77.4 $\pm$ 1.0
HiMol (Zang et al., 2023) 71.3 $\pm$ 0.6 76.0 $\pm$ 0.2 - 62.5 $\pm$ 0.3 70.6 $\pm$ 2.1 84.6 $\pm$ 0.2 - -
Mole-BERT (Xia et al., 2023) 71.9 $\pm$ 1.6 76.8 $\pm$ 0.5 62.8 $\pm$ 1.1 62.8 $\pm$ 1.1 78.9 $\pm$ 3.0 80.8 $\pm$ 1.4 78.6 $\pm$ 1.8 78.2 $\pm$ 0.8
MolCLR (Wang et al., 2022b) 73.3 $\pm$ 1.0 74.1 $\pm$ 5.3 - 61.2 $\pm$ 3.6 89.8 $\pm$ 2.7 82.8 $\pm$ 0.7 78.9 $\pm$ 2.3 77.4 $\pm$ 0.6
GLAD (Nguyen et al., 2024b) 80.4 $\pm$ 1.5 - - 64.7 $\pm$ 1.8 87.3 $\pm$ 1.2 85.7 $\pm$ 0.9 - -
N-GRAM (Hu et al., 2019) 70.8 $\pm$ 1.5 78.7 $\pm$ 0.4 66.5 $\pm$ 0.3 62.7 $\pm$ 0.8 72.6 $\pm$ 1.5 84.5 $\pm$ 0.7 81.3 $\pm$ 2.1 79.9 $\pm$ 0.7
GROVER (Rong et al., 2020) 86.8 $\pm$ 2.2 80.3 $\pm$ 2.0 65.3 $\pm$ 0.5 61.2 $\pm$ 2.5 70.3 $\pm$ 13.7 82.4 $\pm$ 3.6 67.3 $\pm$ 1.8 68.2 $\pm$ 1.1
GraphMVP (Liu et al., 2021) 72.4 $\pm$ 1.6 75.9 $\pm$ 0.5 63.1 $\pm$ 0.4 63.1 $\pm$ 0.4 79.1 $\pm$ 2.8 81.2 $\pm$ 0.9 77.7 $\pm$ 0.6 77.0 $\pm$ 1.2
GEM (Fang et al., 2022) 88.8 $\pm$ 0.4 78.1 $\pm$ 0.4 69.2 $\pm$ 0.4 63.2 $\pm$ 1.5 90.3 $\pm$ 0.7 87.9 $\pm$ 1.1 75.3 $\pm$ 1.5 81.3 $\pm$ 0.3
UniMol (Zhou et al., 2023) 72.9 $\pm$ 0.6 79.6 $\pm$ 0.5 69.6 $\pm$ 0.1 65.9 $\pm$ 1.3 91.9 $\pm$ 1.8 85.7 $\pm$ 0.2 82.1 $\pm$ 1.3 82.8 $\pm$ 0.3
FARM (Ours) 93.3 $\pm$ 0.2 80.8 $\pm$ 1.1 69.9 $\pm$ 0.5 65.9 $\pm$ 0.7 82.2 $\pm$ 0.7 89.6 $\pm$ 0.4 82.7 $\pm$ 2.1 83.5 $\pm$ 0.5
Table 3: Performance comparison of FARM and baseline models on MoleculeNet regression tasks. Performance is evaluated by RMSE and MAE.
Physical chemistry Quantum mechanics
Dataset ESOL Freesolv Lipophilicity QM8 QM9
#tasks 1 1 1 12 3
#samples 1,128 642 4,200 21,786 133,885
Evaluation Metric RMSE ($\downarrow$) MAE ($\downarrow$)
HimGNN (Han et al., 2023) 0.870 $\pm$ 0.154 1.921 $\pm$ 0.474 0.632 $\pm$ 0.016 - -
FG-BERT (Li et al., 2023) 0.944 $\pm$ 0.025 - 0.655 $\pm$ 0.009 - -
Mole-BERT (Xia et al., 2023) 1.015 $\pm$ 0.003 - 0.676 $\pm$ 0.002 - -
N-GRAM (Hu et al., 2019) 1.100 $\pm$ 0.030 2.510 $\pm$ 0.191 0.880 $\pm$ 0.121 0.0320 $\pm$ 0.003 0.00964 $\pm$ 0.00031
GROVER (Rong et al., 2020) 1.423 $\pm$ 0.288 2.947 $\pm$ 0.615 0.823 $\pm$ 0.010 0.0182 $\pm$ 0.001 0.00719 $\pm$ 0.00208
GEM (Fang et al., 2022) 0.813 $\pm$ 0.028 1.748 $\pm$ 0.114 0.674 $\pm$ 0.022 0.0163 $\pm$ 0.001 0.00562 $\pm$ 0.00007
MolCLR (Wang et al., 2022b) 1.113 $\pm$ 0.023 2.301 $\pm$ 0.247 0.789 $\pm$ 0.009 0.0185 $\pm$ 0.013 0.00480 $\pm$ 0.00003
UniMol (Zhou et al., 2023) 0.788 $\pm$ 0.029 1.480 $\pm$ 0.048 0.603 $\pm$ 0.010 0.0156 $\pm$ 0.001 0.00467 $\pm$ 0.00004
FARM (Ours) 0.761 $\pm$ 0.031 1.097 $\pm$ 0.033 0.778 $\pm$ 0.005 0.0146 $\pm$ 0.001 0.00456 $\pm$ 0.00001
@@ 5 Experimental Results
Data and Splits   We consider 13 benchmark tasks in the MoleculeNet dataset (Wu et al., 2018). Following previous works, the data is split using a scaffold split into training, validation, and test sets with an 8:1:1 ratio, ensuring fair and consistent evaluation across all models.
@@
Evaluation Metrics.   We use ROC-AUC as the evaluation metric for classification tasks due to the high imbalance in some datasets. For physical chemistry tasks (ESOL, Freesolv, and Lipophilicity), we use RMSE, and for quantum mechanics tasks (QM8 and QM9), we use MAE, following previous works. For each downstream task, we split the data using three random seeds, train the models, and report the average and standard deviation of the results.

### v4

@@ 3.2 Atom-Level Representation Learning
We employ a masked language model architecture that takes FG-enhanced SMILES as input, using atom-level tokenization and leveraging masked atom prediction as a self-supervised task to train the model. Specifically, we adopt BERT model (Devlin, 2018) with the following loss:
@@
where $\mathcal{M}$ denotes the set of indices corresponding to the masked tokens, $x_{i}$ is the original token at position $i$, and $x_{\setminus i}$ represents the input sequence with the token at position $i$ masked. To evaluate the impact of masking, we conduct experiments with varying masking percentages and test the model’s performance on six downstream tasks. The results indicate that a masking percentage of 35% yields the highest average performance across tasks. Detailed results and information about the training process are provided in Appendix D.1.
@@
Table 1: Evaluation of the FG-enhanced SMILES lexicon size across various databases. “Sequence length” reports the (min; max) number of tokens per molecule. “Vocab size” reports the number of unique tokens included in the lexicon at two minimum frequency thresholds.
Database #Molecules #Atom types Sequence length (min; max) Vocab size (freq $\geq$ 1) Vocab size (freq $\geq$ 5)
ZINC15 3M 10 (5; 63) 1,151 1,089
ChEMBL25 1.8M 35 (1; 867) 15,269 10,016
Collected dataset 20M 46 (1; 867) 22,364 14,741
@@ 3.3 Molecular Structure Learning
To learn the structure of molecules, we employ a two-step process. First, we derive FG embeddings from the FG knowledge graph. Next, we use these embeddings as inputs for a GNN link prediction model to predict interactions between FGs. This approach effectively captures the relationships among FGs, both in terms of their structural characteristics and interactions, thereby facilitating the implicit learning of molecular structure.
@@
where $E^{+}$ denotes the set of positive triples, $E^{-}$ denotes the set of negative triples, and $\gamma$ represents the margin. Optimizing this loss function will maximize the score of positive triples $(\mathbf{h},\mathbf{r},\mathbf{t})$, while minimizing the score of negative triples $(\mathbf{h^{\prime}},\mathbf{r},\mathbf{t^{\prime}})$, with a margin $\gamma$ separating the two. This is achieved through margin-based ranking loss, where the function $f(\mathbf{h},\mathbf{r},\mathbf{t})$ evaluates the plausibility of the triples. The goal is to ensure that the score for positive triples is higher than that for negative triples by at least the margin $\gamma$, thus pushing positive triples closer and negative triples farther apart in the embedding space. The detailed implementation of ComplEx is described in Appendix D.2.
@@
Table 2: Performance comparison of FARM and baseline models on MoleculeNet classification and regression tasks. Classification tasks are evaluated by ROC-AUC (%) ($\uparrow$) and regression tasks by RMSE ($\downarrow$) and MAE ($\downarrow$). The first 11 models incorporate functional group (FG) information. SMI-TED reports QM9 in different units and is not directly comparable.
Classification Regression (RMSE) Regression (MAE)
Physiology Biophysics Physical chemistry Quantum mechanics
Dataset BBBP Tox21 ToxCast SIDER ClinTox BACE MUV HIV ESOL FreeSolv Lipo QM8 QM9
#tasks 1 12 617 27 2 1 17 1 1 1 1 12 3
#samples 2,039 7,831 8,575 1,427 1,478 1,513 93,807 41,127 1,128 642 4,200 21,786 133,885
MICRO 84.4±1.1 77.0±0.8 65.2±0.8 56.7±0.9 77.0±2.0 77.2±2.0 - 75.1±1.1 - - - - -
MGSSL 69.7±0.9 76.5±0.3 64.1±0.7 61.8±0.8 80.7±2.1 79.1±0.9 78.7±1.5 78.8±1.2 - - - - -
MoleOOD 71.0±0.8 - - 63.4±0.7 - 84.3±1.1 - 79.4±0.5 - - - - -
MCM 90.0±3.1 80.2±1.5 - 62.7±2.8 65.5±14 82.0±5.5 - - - - - - -
HimGNN 92.8±2.7 80.7±1.7 - 64.2±2.3 91.7±3.0 85.6±3.4 - - 0.870±0.154 1.921±0.474 0.632±0.016 - -
FG-BERT 70.2±0.9 78.4±0.8 63.3±0.8 64.0±0.7 83.2±1.6 84.5±1.5 75.3±2.4 77.4±1.0 0.944±0.025 - 0.655±0.009 - -
HiMol (S) 71.3±0.6 76.0±0.2 - 62.5±0.3 70.6±2.1 84.6±0.2 - - - - - - -
HiMol (L) 73.2±0.8 76.2±0.3 - 61.3±0.5 80.8±1.4 84.3±0.3 - - - - - - -
Mole-BERT 71.9±1.6 76.8±0.5 62.8±1.1 62.8±1.1 78.9±3.0 80.8±1.4 78.6±1.8 78.2±0.8 1.015±0.003 - 0.676±0.002 - -
MolCLR 73.6±0.5 79.8±0.7 - 68.0±1.1 93.2±1.7 89.0±0.3 78.9±2.3 88.6±2.2 1.113±0.023 2.301±0.247 0.789±0.009 0.0185±0.013 0.00480±0.00003
SCAGE 73.4±1.1 79.4±1.2 69.3±0.5 66.0±1.2 92.7±0.9 85.4±1.2 - - 0.723±0.041 1.688±0.080 0.654±0.009 - -
GLAD 80.4±1.5 - - 64.7±1.8 87.3±1.2 85.7±0.9 - - - - - - -
N-GRAM 70.8±1.5 78.7±0.4 66.5±0.3 62.7±0.8 72.6±1.5 84.5±0.7 81.3±2.1 79.9±0.7 1.100±0.030 2.510±0.191 0.880±0.121 0.0320±0.003 0.00964±0.00031
GROVER 86.8±2.2 80.3±2.0 65.3±0.5 61.2±2.5 70.3±13.7 82.4±3.6 67.3±1.8 68.2±1.1 1.423±0.288 2.947±0.615 0.823±0.010 0.0182±0.001 0.00719±0.00208
GraphMVP 72.4±1.6 75.9±0.5 63.1±0.4 63.1±0.4 79.1±2.8 81.2±0.9 77.7±0.6 77.0±1.2 - - - - -
GEM 88.8±0.4 78.1±0.4 69.2±0.4 63.2±1.5 90.3±0.7 87.9±1.1 75.3±1.5 81.3±0.3 0.813±0.028 1.748±0.114 0.674±0.022 0.0163±0.001 0.00562±0.00007
UniMol 72.9±0.6 79.6±0.5 69.6±0.1 65.9±1.3 91.9±1.8 85.7±0.2 82.1±1.3 82.8±0.3 0.788±0.029 1.480±0.048 0.603±0.010 0.0156±0.001 0.00467±0.00004
SMI-TED289M (Frozen) 91.5±0.5 81.5±0.5 - 66.0±0.9 93.5±0.9 85.6±0.9 - 80.5±1.3 0.705±0.034 1.668±0.062 0.650±0.012 0.0179±0.0004 -
SMI-TED289M (FT) 92.3±0.6 81.9±1.4 - 65.7±0.5 94.3±1.8 88.2±0.5 - 76.9±0.9 0.611±0.010 1.223±0.003 0.552±0.019 0.0095±0.0001 -
FARM (Ours) 93.3±0.2 80.8±1.1 69.9±0.5 65.9±0.7 82.2±0.7 89.6±0.4 82.7±2.1 83.5±0.5 0.761±0.031 1.097±0.033 0.778±0.005 0.0146±0.001 0.00456±0.00001
Link prediction. For link prediction using a graph convolutional network (GCN), the FG graph is used as input. We employ embeddings from the FG knowledge graph as node features for the GCN. During training, node embeddings are computed through graph convolution:
$\mathbf{h}_{i}^{\prime}=\text{ReLU}\left(\mathbf{W}\cdot\frac{1}{|\tilde{\mathcal{N}}(i)|}\sum_{j\in\tilde{\mathcal{N}}(i)}\mathbf{h}_{j}\right)$
where $\mathbf{h}_{i}^{\prime}$ is the updated embedding for node $i$, computed by averaging the embeddings $\mathbf{h}_{j}$ of its neighbors $\tilde{\mathcal{N}}(i)$, including itself due to the added self-loop, applying the weight matrix $\mathbf{W}$, and passing through a ReLU activation function.
The score estimates the probability of connections between nodes is computed with a multi-layer perceptron (MLP) and a sigmoid function: $p_{ij}=\sigma(\text{MLP}(\mathbf{h}_{i}\oplus\mathbf{h}_{j}))$, where $\oplus$ indicates the concatenation operation.
We optimize the model using binary cross-entropy loss defined as: We then sample positive edges $E^{+}$ and negative edges $E^{-}$, and optimize the model using binary cross-entropy loss defined as:
@@ 3.4 Molecular Structure Integration via Contrastive Learning
To integrate atom-level representations with molecular structure information, we employ contrastive learning to align FG-enhanced SMILES representations with FG graph embeddings. This method captures both the atom-level and topological aspects of molecular structures, allowing the model to develop a unified representation that integrates chemical context with overall molecular architecture.
@@
Specifically, given a positive pair $(\mathbf{h}_{\text{MLM}},\mathbf{h}_{\text{pos}})$, where $\mathbf{h}_{\text{MLM}}$ is the atom-level representation derived from a pretrained BERT model and $\mathbf{h}_{\text{pos}}$ is the corresponding structure representation from a graph neural network (GNN), and a negative pair $(\mathbf{h}_{\text{MLM}},\mathbf{h}_{\text{neg}})$, where $\mathbf{h}_{\text{neg}}$ is a augmented FG-graph, the contrastive loss can be written as:
$\mathcal{L}_{\text{CL}}=\frac{1}{N}\sum_{i=1}^{N}\frac{1}{K}\sum_{j=1}^{K}\max\left(0,\ \gamma-\cos\left(\mathbf{h}_{\text{MLM}}^{(i)},\mathbf{h}_{\text{pos}}^{(i)}\right)+\cos\left(\mathbf{h}_{\text{MLM}}^{(i)},\mathbf{h}_{\text{neg}}^{(j)}\right)\right)$
@@
The effect of the contrastive learning objective can be observed in Figure 4(b) and Figure 8. The learned embedding space exhibits meaningful chemical relationships between molecules. The learned embedding space exhibits meaningful chemical relationships between molecules. In particular, similar to the parallel relationships observed in word embeddings in natural language processing, such as king : queen being analogous to man : woman or husband : wife, substituting one functional group with another produces consistent directional shifts in the molecular embedding space across different molecules. This suggests that functional group substitutions correspond to approximately parallel transformations in the learned representation space, indicating that the model captures systematic chemical relationships between functional groups and their structural contexts.
@@ 4 Pre-training Data Collection and Diversity Assessment
Table 3: Performance of FARM on the photostability dataset for quantum mechanics property prediction. $R^{2}$ scores are reported with standard deviation over 10 runs.
Class $\mathbf{R^{2}}$ Class $\mathbf{R^{2}}$ Class $\mathbf{R^{2}}$ Class $\mathbf{R^{2}}$
HOMOm1 (eV) $0.927\pm 0.015$ LUMOp1 (eV) $0.922\pm 0.028$ T3 $0.900\pm 0.010$ S3 $0.911\pm 0.008$
HOMO (eV) $0.895\pm 0.030$ DipoleMoment (D) $0.755\pm 0.053$ T4 $0.921\pm 0.020$ S4 $0.919\pm 0.008$
LUMO (eV) $0.928\pm 0.009$ LogP $0.995\pm 0.006$ T5 $0.905\pm 0.032$ S5 $0.924\pm 0.014$
PrimeExcite (eV) $0.877\pm 0.049$ T1 $0.929\pm 0.006$ S1 $0.936\pm 0.016$ O1 $0.839\pm 0.034$
PrimeExcite (osc) $0.837\pm 0.024$ T2 $0.925\pm 0.018$ S2 $0.948\pm 0.016$ TDOS2.0 $0.868\pm 0.099$
TDOS2.5 $0.877\pm 0.004$ TDOS3.0 $0.815\pm 0.045$ TDOS3.5 $0.895\pm 0.033$ TDOS4.0 $0.919\pm 0.031$
TDOS4.5 $0.938\pm 0.018$
@@ 5.1 Data, Splits and Evaluation Metrics
MoleculeNet We consider 13 benchmark tasks in the MoleculeNet dataset (Wu et al., 2018). Following previous works, the data is split using a scaffold split into training, validation, and test sets with an 8:1:1 ratio, ensuring fair and consistent evaluation across all models. For all downstream tasks, FARM is used as a frozen feature extractor paired with a GRU prediction head; full finetuning details are provided in Appendix D.5.
@@
We use ROC-AUC for classification tasks, RMSE for physical chemistry tasks (ESOL, FreeSolv, and Lipophilicity), and MAE for quantum mechanics tasks (QM8 and QM9), following previous works. For QM9, we report the average MAE across three targets: HOMO energy, LUMO energy, and HOMO-LUMO gap. For each task, we run three random seeds and report the average and standard deviation.
D-B-A Oligomer Photostability Dataset  (Angello et al., 2024) This dataset comprises a chemical library of light-harvesting donor–bridge–acceptor (D–B–A) oligomers, constructed using a modular approach with 22 donor–bridge blocks and 100 acceptor blocks. This design results in a theoretical chemical space of 2,200 unique molecules, which were synthesized and experimentally characterized to study photostability in solution. We adopt a hard-case scaffold split: 3 donor–bridge blocks and 15 acceptor blocks are held out to create the test set, including all molecules containing any of the held-out donors or acceptors, ensuring that no molecules in the training or validation sets share the same donor or acceptor structures. A similar procedure is used to construct the validation set, and the process is repeated for ten random seeds to ensure robust evaluation across multiple train/validation/test splits.
The tasks include the prediction of frontier molecular orbital energies (HOMO, LUMO, HOMO-1, LUMO+1), dipole moment, logarithm of the partition coefficient (LogP), excitation energies (PrimeExcite), oscillator strengths (PrimeExcite (osc)), triplet and singlet excitation energies (T1–T5, S1–S5), and time-dependent density of states (TDOS 2.0–4.5). We use R2 as the evaluation metric to quantify the accuracy of predicted properties against experimental or high-level computational values. Models are trained on the training set and validated on the held-out validation set, with performance reported on the test set across all ten random scaffold splits to ensure robustness and assess generalization to unseen donor or acceptor blocks.
5.2 Baselines